OpenStudy (anonymous):

(Consumer Math) You make an initial investment of $25 at a 3.25% APR that is compounded quarterly for 10 years. You make no other contributions. How much is the interest at the end of 10 years?

OpenStudy (mathstudent55):

Use the compound interest formula for future value.

OpenStudy (anonymous):

Ahh thank you for answering @mathstudent55 have you done these problems before? I just looked up the compound interest formula, is this the right one? A=P(1+r/n)^n*t

OpenStudy (mathstudent55):

Yes, that's it. \(A = P\left(1 + \dfrac{r}{n}\right)^{nt} \) Use P = principal amount = $25 r = interest rate as a decimal = 0.0325 n = number of times interest is compounded in a year = 4 t = time in years = 10

OpenStudy (anonymous):

Thank you Honorary Professor of Mathematics, medal for you :)